1/2(6x+5)-3/4(4x-5)
Answer by oscargut(2103)

1/2(6x+5)-3/4(4x-5)
Solution:
1/2(6x+5)-3/4(4x-5) =
3x+5/2-3x+15/4 =
5/2 +15/4 =
10/4+15/4 =
25/4
Answer: 25/4
You can visit my page: http://maths2012.webs.com
Or you can send me more problems to my e-mail: mthman@gmail.com